### v1

# PAVLM: Advancing Point Cloud based Affordance Understanding Via Vision-Language Model

Abstract
Affordance understanding, the task of identifying actionable regions on 3D objects, plays a vital role in allowing robotic systems to engage with and operate within the physical world. Although Visual Language Models (VLMs) have excelled in high-level reasoning and long-horizon planning for robotic manipulation, they still fall short in grasping the nuanced physical properties required for effective human-robot interaction. In this paper, we introduce PAVLM (Point cloud Affordance Vision-Language Model), an innovative framework that utilizes the extensive multimodal knowledge embedded in pre-trained language models to enhance 3D affordance understanding of point cloud. PAVLM integrates a geometric-guided propagation module with hidden embeddings from large language models (LLMs) to enrich visual semantics. On the language side, we prompt Llama-3.1 models to generate refined context-aware text, augmenting the instructional input with deeper semantic cues. Experimental results on the 3D-AffordanceNet benchmark demonstrate that PAVLM outperforms baseline methods for both full and partial point clouds, particularly excelling in its generalization to novel open-world affordance tasks of 3D objects. For more information, visit our project site: pavlm-source.github.io.
I INTRODUCTION
Affordance understanding is a critical challenge in computer vision, as it is also fundamental to enabling robots to interact effectively with their environment. This understanding is crucial for a wide range of robotic tasks, including human-object interaction [1, 2], spatial localization [3], and most importantly, object manipulation [4]. Traditional methods often rely on supervised learning from extensive human annotation, which can be time-consuming and limited to specific scenarios [5]. Moreover, these approaches struggle to generalize to new objects and unseen environments, a critical limitation in robotic manipulation tasks where flexibility and adaptability are essential [6]. For robotic manipulation, affordance understanding involves more than recognizing an object’s 3D geometry, which requires understanding the object’s functional properties, specifically how it can be grasped, manipulated, or used in different contexts. A holistic concept of affordance understanding in real world environments should also take into account the embodiment of the robot, the configuration of the object in its surroundings, and the intended actions of the agent [7]. In this work, we mainly focus on the affordance understanding of object point clouds in both full shape and partial shape, which potentially serve as a solution for real-world embodied manipulation.
Despite recent advancements in Visual Language Models (VLMs) from popular open-source models such as LLaVA [8], Blip-2 [9], to recent open-source SOTA models like Qwen-VL [10], Phi-3.5-vision [11], Pixtral [12], and many others. These VLM models, while demonstrating strong visual understanding and reasoning capabilities, remain largely confined to 2D visual tasks involving RGB images. They lack the geometric depth required for effective manipulation in 3D affordance-based tasks. Qian et al. [13] were the pioneers in proposing the use of large language models (LLMs) to predict 2D affordance maps by inputting RGB images alongside pseudo-depth information, achieving impressive grounding performance. However, the absence of true depth data limits the robot’s ability to interact effectively in 3D space, where a detailed understanding of the physical structure of an object and its possibilities of interaction is critical [14]. Consequently, while VLMs demonstrate strong abilities in high-level reasoning and following instructions, their application to robotic manipulation remains limited due to their shallow understanding of the physical world.
Figure 1: Comparison of the overview of (a) PointCLIP series and (b) our PAVLM. PointCLIP series methods rely on a pre-trained VLM, such as CLIP, to process language instructions and multi-view depth maps generated from the point cloud through a series of operations like quantization, densification, smoothing, and squeezing. These complex projection-based steps are limited by the number of views, restricting the model’s capacity for comprehensive open-world point cloud understanding. In contrast, our method eliminates the need for these projection steps by employing a geometric-guided point encoder, which directly extracts rich semantic features from the point cloud. Additionally, we fine-tune only a small set of LLM embeddings, maintaining a robust and detailed spatial representation. This approach enhances the understanding of the point cloud from an affordance perspective, providing a more efficient and scalable solution for 3D object interaction.
To overcome these limitations, some researchers have proposed the use of VLMs to process point cloud data by projecting it onto depth maps, as demonstrated by PointCLIP [15] and PointCLIP V2 [16], shown in Figure 1. However, these methods involve multiple projection steps, such as quantization, densification, smoothing, and squeezing, which inevitably lead to a loss of geometric and spatial information. In this paper, we introduce a novel framework that directly extracts semantic features from raw point-cloud data. By moving from 2D representations to 3D point clouds, our model preserves the richness of geometric details and ensures a more comprehensive and accurate affordance understanding. Specifically, our proposed point cloud affordance Vision-Language Model (PAVLM) integrates a geometric-guided propagation module, with the semantic reasoning and world knowledge embedded in large language models (LLMs), such as the Llama series [17, 18]. We also prompt Llama-3.1 models to generate refined context-aware text, augmenting instructional input with deeper semantic cues. Experimental results on the 3D-AffordanceNet benchmark demonstrate that our approach significantly improves performance on open-vocabulary affordance learning tasks and shows strong generalization capabilities, particularly in scenarios involving novel objects and unseen environments. By incorporating world knowledge and 3D geometry, our model generalizes more effectively to novel objects, enabling a more intuitive and effective robotic interaction with the 3D world.
The main contributions of this paper can be summarized as:
  • We propose a unified architecture that integrates 3D point cloud data with an open-source language model for enhanced affordance understanding.
  • We introduce a comprehensive augmented data pipeline for both 3D point cloud objects and corresponding question-answer pairs for each training sample.
  • Our approach outperforms state-of-the-art methods on the 3D-AffordanceNet benchmark, demonstrating a stronger generalization to novel object categories.
II RELATED WORK
Supervised Learning for Affordance Understanding As affordance understanding in the robotic field, it can be simply regarded as object detection and semantic segmentation tasks as a labeling process [19]. However, when a robotic agent interacts with objects, it must recognize them not only by their physical characteristics but also by the actions they enable. Thus, a robust understanding of affordances should improve a robot’s ability to perform appropriate actions tailored to diverse tasks and human requirements. With the rise of deep learning, several methods have emerged that leverage neural networks for learning visual affordance. Notably, [20] introduced an encoder-decoder architecture designed to predict affordance labels from RGB images. Building on this, [21] further advanced affordance detection and segmentation in real, unlabeled data by learning from synthetic RGB datasets. To incorporate richer geometric information, benchmarks such as 3D AffordanceNet [22] and PartAfford [23] were developed, enabling more sophisticated 3D affordance reasoning. These datasets provide dense per-point affordance annotations, illustrating the spatial distribution of interaction regions in 3D space. Moreover, recent advances have inspired context-aware affordance applications such as task-oriented grasp estimation [24] and human hand pose detection [2]. Despite these advancements, most prior work on visual affordance relies on supervised or self-supervised learning, limiting the model’s ability to generalize to novel objects or affordance labels. Our proposed PAVLM overcomes these limitations by harnessing the zero-shot capabilities of pretrained vision language models. By utilizing 3D point cloud data in conjunction with pre-trained language models, our approach significantly extends generalization to unseen object categories.
Pretraining from 3D Point Clouds To effectively obtain feature representations from point clouds, PointNet++ [25] introduced a hierarchical approach to progressive group neighbouring points and learned local features at multiple scales. Drawing inspiration from the concept of mask-based pretraining, Point-MAE [26] and Point-FEMAE [27] introduced a masked auto-encoding scheme to prevent leakage of location information and uneven information density of the point cloud feature. They significantly boost performance in tasks like point cloud reconstruction tasks. Inspired by the success of BERT [28] in natural language processing, Point-BERT [29] adopted the transformer-based Masked Point Modeling (MPM) scheme for point clouds. This method randomly masks out patches of point cloud data and feeds the masked input into a transformer backbone for pre-training. Although these approaches improved point cloud understanding, experiments demonstrated its limited capacity to generalize across new tasks and domains.
The success of transformers in large language models (LLMs) has opened the door to incorporating point cloud features into LLMs. Point-Bind LLM [30], based on ImageBind [31] work, connects six different modalities, including text, images, audio, video, IMU, and depth data, achieving a more comprehensive multimodal understanding. To further enhance point cloud embeddings, ImageBind-LLM [32] provided instruction-based fine-tuning after alignment, enabling the LLM to better interpret inputs and generate more accurate outputs for specific tasks. This process leverages the LLaMA-Adapter [33], introducing an adapter module to fine-tune LLMs for specialized tasks like point cloud processing. Our approach adopts a similar pretraining strategy to leverage open-world knowledge from the pretrained multimodal LLMs, but specifically adapts it to a robot-oriented setting focused on point cloud-based affordance understanding.
Figure 2: The pipeline of our PAVLM. To encode point cloud data directly, we devise a geometric-guided propagation module to obtain point features. On the language side, we enhance textual instructions by prompting Llama-3.1 to generate richer and more detailed guidance. Both visual and textual embeddings are processed through the 3D Image-Bind approach to achieve feature alignment, and the concatenated features are fed into a multi-modal LLM (Llama-2) for mask label generation. Finally, the per-point feature embeddings, derived from the geometric-guided propagation module, are multiplied by the embedding corresponding to the <mask label> token, which is then input into a 3D affordance decoder to generate the corresponding affordance map.
Visual Language Model for Manipulation The emergence of LLM-based intelligent agents with visual understanding, such as LLAVA [34] and ImageBind-LLM [32], has significantly advanced the field of Embodied AI (E-AI) [35]. Visual Language Models (VLMs) have been particularly effective in enabling high-level, long-horizon planning for robotic tasks, leveraging the reasoning capabilities of LLMs. For example, Voxposer [36] combines VLMs with the code-as-policy approach [37] to generate Python code for robotic manipulation tasks. However, this approach is somewhat simplistic, often focusing on locating an object’s midpoint without considering critical affordances, such as grasping a long object like a knife by the handle instead of the middle.
To address this limitation, Copa [38] introduced the Segment Anything Model (SAM), which allows for more precise selection of grasping points, enhancing the generalization abilities of Voxposer. However, since SAM operates only in 2D, it limits planning for robotic manipulation in 3D space. While 2D affordance maps are commonly used, AffordanceLLM [13] was the first to integrate a 2D affordance map with an LLM to improve the agent’s reasoning abilities. Despite this, the gap between 2D affordance reasoning and real-world 3D manipulation remains a challenge. To bridge this gap, we propose the framework of PAVLM to achieve 3D object affordance generation, potentially enabling more effective robotic manipulation in real-world environments.
III The proposed PAVLM Framework
In this section, we present our approach, which takes as input a point cloud $P$ and a textual instruction $T$ to produce a 3D affordance map guided by a <mask label>. The mask label is used to integrate with point cloud embeddings to accurately generate the affordance map for the point cloud.
More detailed, the process begins with the geometric-guided point encoder, which extracts geometric features from the input point cloud. These features are then infused into the point cloud embedding to enrich the representation. Next, we utilize the Llama-3.1 LLM model through prompting to enhance the reasoning capabilities of the system via data augmentation, generating more diverse context-aware question-answer pairs. Afterward, we adopt a pre-trained 3D Image-Bind model to align the point cloud embedding with the textual tokens, ensuring multimodal consistency between the 3D data and the instruction. Following this, the fine-tuned Llama-2 model generates the affordance mask labels, which are integrated with the point cloud embeddings. Finally, these embeddings are fed into a 3D affordance decoder, which produces the final affordance map with high precision. A detailed pipeline of our framework is shown in Figure 2.
Figure 3: Architecture of the proposed Geometric-guided Point Encoder in our PAVLM pipeline.
III-A Geometric-guided Point Encoder
As illustrated in Figure 3, we propose a geometric-guided point encoder to generate the final point cloud embeddings. This encoder is composed of two key submodules: (a) the geometric information extraction module and (b) the feature propagation module. In the first module, we begin by dividing the input point cloud $P\in\mathbb{R}^{2048\times 3}$ into $k$ point patches. Inspired by Point-Bert [29], we applied convolutional layers to extract features from these patches, resulting in point-wise features $\{f_{i}\}_{i=1}^{k}$. These features are then combined with positional embeddings $\{p_{i}\}_{i=1}^{k}$ and fed into a series of transformer blocks via a cross-attention mechanism. This process yields refined geometric features $\{g_{i}\}_{i=1}^{k}$. We adopt a structure comprising 3-layer transformer blocks, and the extraction of geometric features can be formulated as follows:
$g_{f}=\mathcal{G}(f_{i},p_{i}),\quad\hat{F}_{j}=F_{j}+g_{j}\cdot\text{softmax}\left(\frac{F_{j}g_{f}^{T}}{\sqrt{D}}\right)g_{f},$ (1)
where $F_{j}$ represents the output feature from the $j$-th transformer layer. To enhance the quality of point embeddings, we incorporate a feature propagation module, as recommended by [39]. Specifically, we process the output features to generate high-quality per-point embeddings. We employ an upsampling technique to propagate the features from the first and second transformer block outputs, whereas Dynamic Graph CNN (DGCNN [40]) propagation techniques are also used to ensure the features are effectively distributed across the point cloud.
III-B Textual Instruction Augmentation
To enhance affordance understanding through the alignment of point cloud data and text, we incorporate 3D-specific descriptions that reflect implicit human-object interaction characteristics as the textual input. Rather than using general action terms like cut, handover, and tool use, commonly seen in previous affordance-based robotic works [24, 41], we adopt a more detailed question-answering approach inspired by AffordanceLLM [13]. In this approach, a seed question is generated, such as ”What part of the <object name> should we interact with to <action name> it?” followed by an answer like ”You can <action name> the area <mask token>.”
Given the powerful descriptive capabilities of large language models (LLMs), we extend this approach by utilizing Llama-3.1 70B [42] to generate richer, 3D-specific augmentation question-answering pairs. Llama-3.1 processes the language command and outputs responses based on its pretrained knowledge, resulting in more detailed and context-aware interactions. An example of the prompt we used for this process is illustrated in Figure 4. For a 3D dataset with $K$ categories, we replace the <mask label> in each command with the corresponding category value and input these commands into Llama-3.1, generating 3D-specific descriptions enriched with category-specific semantics.
Figure 4: The augmentation prompt template of our textual instructions to enhance the diversity of question and answer pairs.
III-C Feature Alignment of Point Cloud and Text
Inspired by Image-Bind [31], it integrates five modalities, including video, depth, IMU, audio, text, and thermal data to achieve cross-model feature learning. Point-Bind [30] extends this idea by constructing a joint embedding space between 3D data and multiple modalities. After training on large-scale point-paired data, Point-Bind effectively aligns point cloud data with images, text, and audio into a unified representation space. By introducing a contrastive loss function, Point-Bind achieves 3D-centric multimodal learning. Building on this idea, we align the extracted feature from the point cloud $P$ and text $T$ with the contrastive batch average loss $L_{c}$ by calculating their Euclidean distance and maintain the minimum margin as 1, shown in Equation  2.
$\begin{split}L_{ca}=\text{Average}(\text{dist}(P,T))\\ =\text{Average}\|P-T+\textit{eps}\cdot e\|_{p},\end{split}$ (2)
where $\textit{eps}\cdot e$ term is used to avoid numerical instability or division by zero in distance calculations.
III-D Masked-based Affordance Decoding
Continually, the aligned feature embeddings generated by 3D Image-Bind are fed into a pretrained multimodal Llama-2 model. In this process, we freeze most of the Llama-2 weight parameters and apply the fine-tuning strategy of the LLaMA-Adapter [33] to predict a special <mask label>. Based on the 3D AffordanceNet benchmark [43], we define 18 categories of affordance mask labels. To encode the output labels, we employ a one-hot encoding scheme, transforming the hidden state of the <mask label> into a query embedding $q$.
Seen in Figure 3, this query embedding $q$ first concatenates with the original point cloud $P$, and then integrated with point cloud embedding $P_{em}$ into the 3D affordance decoder, which consists of two convolutional layers and a batch normalization layer. Its primary function is to reduce the dimensionality of the multimodal feature embeddings and produce a point-wise affordance map $\mathbf{M}\in\mathbb{R}^{2048\times 1}$ in the point cloud, where each point $M_{i}\in[0,1]$ represents the affordance score for point $i$ in the point cloud. The above process can be mathematically expressed as:
$\mathbf{M}=\text{Decoder}(P_{em},(P\oplus q)),$ (3)
$\mathbf{M}=\{M_{i}\mid M_{i}\in[0,1],\,i=1,2,\dots,2048\}.$ (4)
III-E Training Objectives
For the comprehensive training of both the geometric-guided encoder and the 3D affordance decoder, we adopt a multi-objective loss function that integrates several components to ensure effective learning. The core affordance loss $L_{aff}$ is composed of a cross-entropy loss $L_{ce}$ and a Dice loss $L_{dice}$, following the practice established in prior work [43]. This combination helps to optimize both the pixel-wise classification and spatial consistency of the affordance predictions. In addition to the affordance loss, we introduce an alignment loss to ensure that the point cloud embeddings are well aligned with the textual modality, as defined in Eq. 2. This alignment is crucial for integrating 3D geometric information into the joint representation space. Furthermore, when fine-tuning the Llama-2 model with LLaMA-Adapter, we introduce an additional query loss $L_{q}$, which is a cross-entropy loss aimed at improving the learning of affordance labels based on the query embeddings generated from the textual prompts. The overall affordance learning loss during the training process is defined as:
$L_{aff}=L_{ce}+\lambda\cdot L_{dice},$ (5)
where the weighting factor $\lambda$ is set to 1. Both the alignment loss $L_{ca}$ and the query loss $L_{q}$ are used independently to fine-tune the pre-trained multimodal model.
IV Experiments
Given a 3D point cloud of an object, our objective is to estimate affordances for both complete and partial shapes. In our experiments, we systematically evaluate the performance of our approach by addressing two key questions: (1) How do specific design choices, such as varying prompts and visual encoders, affect overall performance? (2) How well does our method generalize in comparison to state-of-the-art techniques?
IV-A Experiment Setup
Metrics. We assess affordance estimation using four metrics: mean Average Precision (mAP), Area Under the ROC Curve (AUC), average Intersection Over Union (aIoU), and mean squared error (MSE). Firstly, we generate a Precision-Recall curve and compute the Average Precision for each affordance. The area under the ROC curve is calculated as the AUC metric. We also compute the mean squared error (MSE) for each affordance class and sum the results across all categories. After binarizing the predictions in a step of 0.01, the aIoU is calculated by progressively increasing the threshold from 0 to 0.99. Aside from MSE, all metrics are further process In an average mathematical operation.
Baselines. We compare our approach with state-of-the-art baseline methods. As shown in Figure 1, we mainly compare our work with PointCLIP [15] and PointCLIP V2 [16], which both achieve great generalization performance on the language-interactive 3d point cloud understanding tasks.
  • PointCLIP [15]: PointCLIP is the first model to bridge the gap between the 2D image-text pair pre-trained model (CLIP) and the irregular structure of 3D point clouds. It introduces a multi-view projection-based method that transform point clouds into multi-view depth maps, which are then processed by CLIP’s visual encoder. It achieves promising results on standard 3D benchmarks such as ModelNet and ScanObjectNN.
  • PointCLIP V2 [16]: PointCLIP V2 is an enhanced framework that integrates CLIP and GPT models for powerful 3D open-world learning. Building on the original PointCLIP, PointCLIP V2 introduces more realistic visual projections and rich 3D-specific text prompts, making it a robust and versatile tool for open-world point cloud understanding.
TABLE I: Ablation study of different text prompts.
[…]
Full-shape (Aug.) Ours 45.7 86.9 16.8 0.45
Full-shape Ours 46.0 87.1 17.1 0.44
Full-shape Object, Action 44.5 85.6 16.0 0.67
Full-shape Action 45.4 87.2 16.7 0.45
Full-shape Hi 11.4 51.9 0.4 0.90
Dataset. We primarily evaluate our methods using the 3D AffordanceNet benchmark [43], which is specifically designed for visual affordance understanding in 3D. The dataset contains 22,949 3D object models across 23 semantic categories and covers 18 affordance categories. During the training process, we employ two dataset-splitting strategies: seen and unseen. For the seen strategy, we shuffle the original dataset and split it into training, validation, and test sets using an 8:1:1 ratio where each of these splits contains all categories. In contrast, the unseen strategy follows the same splitting ratio but introduces specific object categories, namely ”mug,” ”knife,” and ”scissors”, which are entirely reserved for the test set. These categories are excluded from the training and validation sets, making this split particularly valuable for evaluating the model’s ability to generalize to novel objects, which is significant for real-world robotic manipulation tasks.
IV-B Ablation Study
We performed several ablation studies to evaluate how various modules of our model contribute to its overall performance on the seen split dataset. Specifically, we examined the effects of varying text prompts and point cloud encoders. Following the evaluation prompts used in previous work [2], we tested four different types of text prompts to analyze their impact on model performance: (1) A simple greeting noise, such as ”Hi”, (2) The action associated with the object, (3) A combination of the object name and action label, and (4) A full question prompt and its augmentation, such as ”Which specific region of the mug should we target to provide drinking?”.
The results, summarized in Table I, indicate that comprehensive question prompts consistently outperform simpler ones across all four metrics. Notice that when the context closely matches the original dataset prompt, better results are produced. Compared to object names, action names combined with point cloud embeddings provide stronger hints toward the result, indicating that object names are less correlated with the outcome than their associated actions. Furthermore, when textual enhancement is applied, the performance metrics are slightly lower than those of the original prompts. This is expected, as the augmented dataset enhances the LLM’s generalization capability. Although the exact sentence text yields better results, the model becomes more adept at handling complex language instructions.
In addition, we compared our geometric-guided point encoder against two widely used point cloud-based network architectures: PointNet++ [25] and DGCNN [40]. As demonstrated in Table II, our proposed geometric-guided point encoder average outperforms them when used as the vision backbone. For instance, in key metrics like mAP and aIoU, our model achieves over 10% higher performance than the best PointNet++ model. This improved performance indicates that our encoder is more effective in extracting meaningful semantic feature embeddings for our affordance understanding task.
TABLE II: Ablation study of different vision encoders.
[…]
Full-shape PointNet++ 47.1 86.2 18.1 0.36
Full-shape DGCNN 31.0 72.9 1.0 0.68
Full-shape Ours 48.5 86.8 17.7 0.36
Partial-shape PointNet++ 25.3 69.8 3.4 0.73
Partial-shape DGCNN 26.8 69.7 0.7 0.67
Partial-shape Ours 40.2 82.2 11.7 0.39
IV-C Evaluation Result
Comparison with Baselines. To further evaluate 3D affordance understanding, we compared the performance of our model with state-of-the-art point cloud understanding approaches, specifically PointCLIP [15] and PointCLIP V2 [16]. Since both baseline methods leverage the frozen pre-trained vision-language model CLIP, we introduced a variant of our model by freezing the pre-trained Image-Bind and Llama-2 modules. The results presented in Table III, indicate that both baseline methods struggle with our 3D affordance learning task. In contrast, our model, including the frozen variant, consistently delivers superior performance across all four metrics, demonstrating its effectiveness on both full-shape and partial-shape point cloud data.
TABLE III: Comparison results with baselines on the seen split of dataset
[…]
Full-shape PointCLIP 7.6 49.9 0.9 0.80
Full-shape PointCLIP V2 7.6 50.0 0.8 0.70
Full-shape Ours (frozen) 46.3 86.9 17.1 0.44
Full-shape Ours 48.5 86.8 17.7 0.36
Partial-shape PointCLIP 9.1 50.1 1.1 0.78
Partial-shape PointCLIP V2 9.2 49.9 1.4 0.66
Partial-shape Ours (frozen) 37.2 64.2 11.4 0.36
Partial-shape Ours 40.2 82.2 11.7 0.39
Generalization to Open-world Categories. To evaluate the generalization capability of our model to open-world object categories, we conducted experiments on the unseen split of the dataset, using the same experimental setup. As presented in Table IV, the performance of all models, across both full shape and partial shape evaluations, shows a noticeable decline compared to the results on the seen split, as reported in Table III. Despite this overall drop, our model significantly outperforms all baseline methods, particularly in terms of the mAP and MSE metrics. This highlights its strong ability to generalize to novel object categories, even in the challenging task of affordance understanding.
IV-D Visualization Result
As illustrated in Fig 5, we present several qualitative results from our model on unseen objects and categories. When given an implicit instruction, such as ”Where on the chair should we use it for sitting?”, our model accurately identifies the corresponding affordance region on the chair. In some categories from the unseen split, like knives, more complex instructions that require extensive reasoning ability, such as ”To use the knife as a tool, which area should we interact with?”, are handled effectively by our model. It accurately predicts the handle centre of the knife. These results demonstrate that our method leverages the reasoning capabilities of LLMs to generate high-quality affordance masks, even for challenged partial point cloud understanding.
V CONCLUSIONS
We have introduced PAVLM, a novel framework for 3D point cloud-based affordance understanding that successfully handles unseen, open-world objects. Our approach leverages the extensive knowledge embedded in Multimodal Large Language Models (MLLMs) to enhance performance, which significantly surpasses state-of-the-art affordance understanding models, excelling in both full-shape and partial-shape point clouds, even under complex language instructions. This enhanced capability is vital for the development of embodied robots that can adeptly grasp and manipulate objects in dynamic and human-robot collaborative environments. In future work, we plan to explore how to adapt the proposed PAVLM framework to further improve its capabilities and extend its applications to integrate with 6-DOF grasp generation algorithms, enabling its application in real-world robotic tasks.
TABLE IV: Comparison results with baselines on the unseen split of dataset
[…]
Full-shape PointCLIP 3.7 49.9 0.5 1.58
Full-shape PointCLIP V2 3.7 49.3 0.4 0.91
Full-shape Ours (frozen) 11.8 60.3 1.5 0.78
Full-shape Ours 12.1 53.9 2.6 0.51
Partial-shape PointCLIP 4.7 50.2 0.6 1.14
Partial-shape PointCLIP V2 4.7 50.3 0.38 0.93
Partial-shape Ours (frozen) 13.8 60.3 2.0 0.60
Partial-shape Ours 22.3 63.2 2.1 0.82
Figure 5: Visualization of affordance prediction results on multiple unseen objects and categories, showing that our model predicts the accurate affordance as different instructions guides from the original partial point cloud.
VI Acknowldegements
This research was funded by the German Research Foundation (DFG) and the National Science Foundation of China (NSFC) in the project Crossmodal Learning, DFG TRR-169/NSFC 62061136001, and Hon Hai Research Institute (HHRI).